# A localized consensus-based sampling algorithm

### v2

@@ 4.1 Consensus-based sampling without Gaussian assumption
The final approximation, from Eq. 25 to Eq. 27, is only justified for Gaussian $\pi$ and will be omitted. In addition, we keep the $\kappa$ and $\beta$ parameters independent. CBS sets $\kappa=\beta=\alpha$ and compensates this with a suitable $\gamma$ to still sample from Gaussian distributions exactly, but in the more challenging non-Gaussian case they need to be selected carefully. The Moreau envelope approximation to $V$ suggests that $\kappa$ should be chosen small, while the weighted-mean approximation to the proximal operator implies that $\beta$ should be large. Finally, we note that the steps leading up to Eq. 25 are not specific to the preconditioner $\mathcal{C}_{\alpha}(\rho_{t})$ and instead allow a general positive definite preconditioner $\mathbf{C}_{t}\coloneqq\mathbf{C}(\rho_{t})$. This results in our novel localized CBS dynamics
$\dot{U}_{t}=-\frac{\gamma}{\kappa}\Bigl{(}U_{t}-\mu_{\beta,\kappa\mathbf{C}_{t}/\beta}(\rho_{t};U_{t})\Bigr{)}+\sqrt{2\mathbf{C}_{t}}\,\dot{W}_{t}.$ (28)
@@
The parameter $\gamma$ could simply be set to $1$, but other choices may be able to partially mitigate the bias that is due to $\beta<\infty$ and ${\kappa>0}$. We study this bias in more detail in subsection 4.3 as a function of $\gamma$, $\kappa$, and the choice of $\mathbf{C}_{t}$. Notably, when $\pi$ is Gaussian and $\mathbf{C}_{t}$ is an unweighted or weighted covariance—choices (ii) and (iii) above—the bias can be removed fully. The value of $\gamma$ that accomplishes this, given by Eqs. 38 and 39, can then be used in Eq. 28 even in non-Gaussian cases as a heuristic.
@@ 4.2 Localized preconditioners
$\dot{U}_{t}=-\frac{\gamma}{\kappa}\Bigl{(}U_{t}-\mu_{\beta,\kappa\mathbf{C}_{U,t}/\beta}(\rho_{t};U_{t})\Bigr{)}+\nabla_{u}\cdot\mathbf{C}_{U,t}+\sqrt{2\mathbf{C}_{U,t}}\,\dot{W}_{t}.$ (29)
@@
    a weighted covariance $\mathbf{C}_{U,t}=\mathcal{C}_{\alpha,\lambda\mathcal{C}(\rho_{t})}(\rho_{t};U_{t})$ with scalars $\alpha\geq 0$ and $\lambda>0$.
Here,
@@
which uses the weighted mean Eq. 23 and weight function Eq. 24. With this preconditioner, the bias of localized CBS can be removed entirely for Gaussian distributions, similarly to the covariances in subsection 4.1, with the $\gamma$ value given by Eq. 40. This weighted covariance is interesting for comparison to the polarized CBS method discussed in subsection 5.2, since $\mathcal{C}_{\alpha,\lambda\mathcal{C}(\rho_{t})}(\rho;u)$ corresponds to the polarized covariance Eq. 54b with a Gaussian kernel [^4] Polarized CBS does not include $\mathcal{C}(\rho_{t})$ in the distance norm, so the correspondence is not perfect. We add this anisotropy in our covariance weights to ensure affine-invariance (see subsection 4.4). $k(u,v)=\exp(-(2\lambda)^{-1}\lVert u-v\rVert_{\mathcal{C}(\rho_{t})}^{2})$. Using such a weighted covariance can incur a significant extra computational cost and, in our experiments, does not seem to be needed for good performance of localized CBS.
@@ 4.3 Analysis for Gaussian distributions
While the localized CBS dynamics Eq. 29 are designed to sample from non-Gaussian distributions, we will now analyze their behavior in detail in the case where the target distribution $\pi$ is Gaussian, i.e., when ${\pi(u)\sim\widehat{\pi}(u)=\exp(-V(u))=\exp(-\frac{1}{2}\lVert u-m\rVert_{\Sigma}^{2})}$ for some $m$ and $\Sigma$. This will result in values for $\gamma$ that we propose to use even in the non-Gaussian case.
@@
Assume that the target distribution $\pi=\mathcal{N}(m,\Sigma)$ and the initial condition $U_{0}\sim\mathcal{N}(m_{0},\Sigma_{0})$ are Gaussian. Then, the localized CBS dynamics Eq. 29 admit a Gaussian solution $U_{t}\sim\mathcal{N}(m_{t},\Sigma_{t})$ with mean and covariance that satisfy
$\dot{m}_{t}$ $=-\frac{\gamma\beta}{\kappa}P_{t}\Sigma^{-1}(m_{t}-m),$ (32a)
$\dot{\Sigma}_{t}$ $=2\Bigl{[}\mathbf{C}_{*,t}-\frac{\gamma}{\kappa}\Sigma_{t}+\frac{\gamma\beta}{2\kappa^{2}}\left(P_{t}\mathbf{C}_{*,t}^{-1}\Sigma_{t}+\Sigma_{t}\mathbf{C}_{*,t}^{-1}P_{t}\right)\Bigr{]},$ (32b)
where
$P_{t}\coloneqq\Bigl{(}\beta\Sigma^{-1}+\Sigma_{t}^{-1}+\frac{\beta}{\kappa}\mathbf{C}_{*,t}^{-1}\Bigr{)}^{-1},$ (33)
if the preconditioner $\mathbf{C}_{*,t}\coloneqq\mathbf{C}(\rho_{t};U_{t})$ does not depend on $U_{t}$ when $\rho_{t}=\mathcal{N}(m_{t},\Sigma_{t})$. This is the case, for example, for the preconditioners considered before:
@@
Assume that $\rho_{t}=\mathcal{N}(m_{t},\Sigma_{t})$. From the condition that $\mathbf{C}_{*,t}$ not depend on $U_{t}$ follows ${\nabla_{u}\cdot\mathbf{C}_{*,t}=0}$. Similarly to [4], in Appendix A we derive the rightmost terms in Eqs. 34b, 34c and 34d and
@@
In this case, Eq. 29 becomes a linear SDE: the drift term depends on $U_{t}$ in an affine way and the diffusion term is $U_{t}$-independent. Following, e.g., [51, Section 6.1], the SDE admits a Gaussian solution with mean and covariance that satisfy
$\dot{m}_{t}$ $=-\frac{\gamma}{\kappa}(I-\frac{\beta}{\kappa}P_{t}\mathbf{C}_{*,t}^{-1})m_{t}+\frac{\gamma}{\kappa}P_{t}\,(\beta\Sigma^{-1}m+\Sigma_{t}^{-1}m_{t}),$ (36)
$=-\frac{\gamma}{\kappa}\Bigl{[}(I-P_{t}\,(P_{t}^{-1}-\beta\Sigma^{-1}))m_{t}-P_{t}\beta\Sigma^{-1}m\Bigr{]}=-\frac{\gamma\beta}{\kappa}P_{t}\Sigma^{-1}\,(m_{t}-m)$
and
$\dot{\Sigma}_{t}=-\frac{\gamma}{\kappa}(I-\frac{\beta}{\kappa}P_{t}\mathbf{C}_{*,t}^{-1})\Sigma_{t}-\frac{\gamma}{\kappa}\Sigma_{t}\,(I-\frac{\beta}{\kappa}\mathbf{C}_{*,t}^{-1}P_{t})+2\mathbf{C}_{*,t},$ (37)
which simplifies into Eq. 32b. ∎
@@
    $\gamma=\left(\frac{1}{\lambda}+\alpha+1\right)^{-1}\kappa+\frac{\beta}{\beta+1}.$ (40)
@@
For $\pi$ to be a stationary distribution of localized CBS, we need ${\partial_{t}\rho_{t}(u)=0}$ when ${\rho_{t}=\pi}$. Since then both $\rho_{t}$ and $\pi$ are Gaussian, we can use Lemma 4.1, fill in an expression for the preconditioner such as Eq. 34, and require $\dot{m}_{t}=0$ and $\dot{\Sigma}_{t}=0$ when $m_{t}=m$ and $\Sigma_{t}=\Sigma$. Clearly, $\dot{m}_{t}$ as given by Eq. 32a is zero irrespective of $\gamma$ and $\mathbf{C}_{*,t}$ when $m_{t}=m$. For $\dot{\Sigma}_{t}$, we have
$\dot{\Sigma}_{t}$ $=2\Bigl{[}\mathbf{C}_{*,t}-\frac{\gamma}{\kappa}\Sigma+\frac{\gamma\beta}{2\kappa^{2}}\left(P_{t}\mathbf{C}_{*,t}^{-1}\Sigma+\Sigma\mathbf{C}_{*,t}^{-1}P_{t}\right)\Bigr{]}.$ (41)
@@
$\dot{\Sigma}_{t}$ $=2\Bigl{[}\mathbf{C}_{*,t}+\frac{\gamma}{\kappa}\Bigl{(}\bigl{(}\kappa\frac{\beta+1}{\beta}\mathbf{C}_{*,t}\Sigma^{-1}+I\bigr{)}^{-1}-I\Bigr{)}\Sigma\Bigr{]}$ (42)
@@
For the weighted covariance, Eq. 34d becomes $\mathbf{C}_{*,t}=(\lambda^{-1}+\alpha+1)\Sigma$ when $\Sigma_{t}=\Sigma$, which means that Eq. 43 is satisfied when $\gamma$ is given by Eq. 40. The other cases correspond to specific choices for $\lambda$ and $\alpha$ and are analogous. ∎
@@ 4.4 Affine-invariance
The SDE Eq. 29 is affine-invariant when the preconditioner is given by $\mathbf{C}(\rho_{t};U_{t})=\mathcal{C}(\rho_{t})$, $\mathbf{C}(\rho_{t};U_{t})=\mathcal{C}_{\alpha}(\rho_{t})$, or $\mathbf{C}(\rho_{t};U_{t})=\mathcal{C}_{\alpha,\lambda\mathcal{C}(\rho_{t})}(\rho_{t};U_{t})$.
@@
[…]
@@ 5 Comparison to alternative methods
This section compares the dynamics Eq. 29 of localized CBS to those of various other McKean–Vlasov-based samplers. A detailed numerical comparison follows in section 7; here, we focus on different aspects of the dynamics and how these may affect the properties of the resulting algorithms.
@@ 5.1 Consensus-based sampling
Let $U_{t}\sim\mathcal{N}(m_{t},\Sigma_{t})$ follow the localized CBS dynamics Eq. 29 preconditioned with the weighted covariance $\mathbf{C}(\rho;u)=\mathcal{C}_{\alpha,\lambda\mathcal{C}(\rho)}(\rho;u)$ and let $U_{t}^{\prime}\sim\mathcal{N}(m_{t}^{\prime},\Sigma_{t}^{\prime})$ follow the CBS dynamics Eq. 15. The target density $\pi=\mathcal{N}(m,\Sigma)$ and initial densities $\mathcal{N}(m_{0},\Sigma_{0})=\mathcal{N}(m_{0}^{\prime},\Sigma_{0}^{\prime})$ are Gaussian.
@@
$\dot{\Sigma}^{\prime}_{t}=-2\alpha^{\prime}\Sigma^{\prime}_{t}\,(\Sigma+\alpha^{\prime}\Sigma^{\prime}_{t})^{-1}(\Sigma^{\prime}_{t}-\Sigma).$ (51)
@@
$\dot{\Sigma}_{t}$ $=2\Bigl{[}\mathbf{C}-2\frac{\alpha^{\prime}}{\alpha(\alpha^{\prime}+1)}\Sigma_{t}+\frac{(\alpha^{\prime})^{2}}{\alpha^{2}(\alpha^{\prime}+1)}\left(P_{t}\mathbf{C}^{-1}\Sigma_{t}+\Sigma_{t}\mathbf{C}^{-1}P_{t}\right)\Bigr{]}$
@@
where we noted that $\gamma=2\frac{\alpha^{\prime}}{\alpha^{\prime}+1}$. This is the same ODE as Eq. 51, but accelerated with a factor $c$; hence, $\Sigma_{t}=\Sigma^{\prime}_{ct}$.
@@
$\dot{m}^{\prime}_{t}=-\alpha^{\prime}\Sigma^{\prime}_{t}\,(\Sigma+\alpha^{\prime}\Sigma^{\prime}_{t})^{-1}(m^{\prime}_{t}-m).$ (52)
@@
$\dot{m}_{t}$ $=-2\alpha^{\prime}cP_{t}\Sigma^{-1}\,(m_{t}-m)=-c\alpha^{\prime}\Sigma_{t}\,(\Sigma+\alpha^{\prime}\Sigma_{t})^{-1}(m_{t}-m)$

### v3

@@ 4.1 Consensus-based sampling without Gaussian assumption
The final approximation, from Eq. 24 to Eq. 26, is only justified for Gaussian $\pi$ and will be omitted. In addition, we keep the $\kappa$ and $\beta$ parameters independent. CBS sets $\kappa=\beta=\alpha$ and compensates this with a suitable $\gamma$ to still sample from Gaussian distributions exactly, but in the more challenging non-Gaussian case they must be selected carefully. The Moreau envelope approximation Eq. 18 to $V$ suggests that $\kappa$ should be chosen small, while the weighted-mean approximation Eq. 22 to the proximal operator implies that $\beta$ should be large. Finally, we note that the steps leading up to Eq. 24 are not specific to the preconditioner $\mathcal{C}_{\alpha}(\rho_{t})$ and allow a general positive definite preconditioner $\mathbf{C}_{t}\coloneqq\mathbf{C}(\rho_{t})$. This results in our novel localized CBS dynamics
$\mathrm{d}U_{t}=-\frac{\gamma}{\kappa}\Bigl(U_{t}-\mu_{\beta,\kappa\mathbf{C}_{t}/\beta}(\rho_{t};U_{t})\Bigr)\,\mathrm{d}t+\sqrt{2\mathbf{C}_{t}}\,\mathrm{d}W_{t}.$ (27)
@@
The parameter $\gamma$ could simply be set to $1$, but other choices may be able to partially mitigate the bias that is due to $\beta<\infty$ and $\kappa>0$}. We study this bias in more detail in subsection 4.3 as a function of $\gamma$, $\kappa$, and the choice of $\mathbf{C}_{t}$. Notably, when $\pi$ is Gaussian and $\mathbf{C}_{t}$ is an unweighted or weighted covariance—choices (ii) and (iii) above—the bias can be removed fully. The value of $\gamma$ that accomplishes this, given by Eqs. 36 and 37, can then be used in Eq. 27 even in non-Gaussian cases as a heuristic.
@@ 4.2 Localized preconditioners
$\mathrm{d}U_{t}=-\frac{\gamma}{\kappa}\Bigl(U_{t}-\mu_{\beta,\kappa\mathbf{C}_{U,t}/\beta}(\rho_{t};U_{t})\Bigr)\,\mathrm{d}t+\nabla_{u}\cdot\mathbf{C}_{U,t}+\sqrt{2\mathbf{C}_{U,t}}\,\mathrm{d}W_{t}.$ (28)
@@
    a weighted covariance $\mathbf{C}_{U,t}=\mathcal{C}_{\alpha,\lambda\mathcal{C}(\rho_{t})}(\rho_{t};U_{t})$ with scalars $\alpha\geq 0$ and $\lambda>0$, where
@@
Equation 30 uses the weighted mean and weight function Eq. 23. With this preconditioner, the bias of localized CBS can be removed for Gaussian $\pi$, similarly to the covariances in subsection 4.1, with the $\gamma$ value given by Eq. 38. This weighted covariance is interesting for comparison to the polarized CBS method discussed in subsection 5.2, since $\mathcal{C}_{\alpha,\lambda\mathcal{C}(\rho_{t})}(\rho;u)$ corresponds to the polarized covariance Eq. 52b with a Gaussian kernel [^4] Polarized CBS does not include $\mathcal{C}(\rho_{t})$ in the distance norm, so the correspondence is not perfect. We add this anisotropy in our covariance weights to ensure affine-invariance (see subsection 4.4). $k(u,v)=\exp(-(2\lambda)^{-1}\lVert u-v\rVert_{\mathcal{C}(\rho_{t})}^{2})$. Using such a weighted covariance can incur a significant extra computational cost and, in our experiments, does not seem necessary for good performance of localized CBS.
@@ 4.3 Analysis for Gaussian distributions
While the localized CBS dynamics Eq. 28 are designed to sample from non-Gaussian distributions, we will now analyze their behavior in the case where the target distribution is Gaussian, i.e., when $\pi(u)\propto\widehat{\pi}(u)=\exp(-\frac{1}{2}\lVert u-m\rVert_{\Sigma}^{2})$} for some $m$ and $\Sigma$. This can be seen as verifying a minimum requirement: we can hardly expect the method to work well for non-Gaussian distributions if it cannot handle Gaussian ones. Moreover, this analysis results in values for $\gamma$ that we will later use even for non-Gaussian $\pi$.
@@
Assume that both the target distribution $\pi=\mathcal{N}(m,\Sigma)$ and the initial condition $U_{0}\sim\mathcal{N}(m_{0},\Sigma_{0})$} are Gaussian. Then, the localized CBS dynamics Eq. 28 admit a Gaussian solution $U_{t}\sim\mathcal{N}(m_{t},\Sigma_{t})$ with mean and covariance that satisfy
$\frac{\mathrm{d}}{\mathrm{d}t}m_{t}$ $=-\frac{\gamma\beta}{\kappa}P_{t}\Sigma^{-1}(m_{t}-m),$ (31a)
$\frac{\mathrm{d}}{\mathrm{d}t}\Sigma_{t}$ $=2\Bigl[\mathbf{C}_{*,t}-\frac{\gamma}{\kappa}\Sigma_{t}+\frac{\gamma\beta}{2\kappa^{2}}\left(P_{t}\mathbf{C}_{*,t}^{-1}\Sigma_{t}+\Sigma_{t}\mathbf{C}_{*,t}^{-1}P_{t}\right)\Bigr],$ (31b)
with $P_{t}\coloneqq(\beta\Sigma^{-1}+\Sigma_{t}^{-1}+\frac{\beta}{\kappa}\mathbf{C}_{*,t}^{-1})^{-1}$, if the preconditioner $\mathbf{C}_{*,t}\coloneqq\mathbf{C}(\rho_{t};U_{t})$ does not depend on $U_{t}$ when $\rho_{t}=\mathcal{N}(m_{t},\Sigma_{t})$. This is the case for the preconditioners considered before:
@@
Assume $\rho_{t}=\mathcal{N}(m_{t},\Sigma_{t})$. From the condition that $\mathbf{C}_{*,t}$ not depend on $U_{t}$ follows that $\nabla_{u}\cdot\mathbf{C}_{*,t}=0$}. Appendix A derives the rightmost terms in Eqs. 32b, 32c and 32d and
@@
In this case, Eq. 28 becomes a linear SDE: the drift term depends on $U_{t}$ in an affine way and the diffusion term is $U_{t}$-independent. Following, e.g., [52, Section 6.1], the SDE admits a Gaussian solution with mean and covariance that satisfy
$\begin{aligned} \frac{\mathrm{d}}{\mathrm{d}t}m_{t}&=-\frac{\gamma}{\kappa}(I-\frac{\beta}{\kappa}P_{t}\mathbf{C}_{*,t}^{-1})m_{t}+\frac{\gamma}{\kappa}P_{t}\,(\beta\Sigma^{-1}m+\Sigma_{t}^{-1}m_{t})\\ &=-\frac{\gamma}{\kappa}\Bigl[(I-P_{t}\,(P_{t}^{-1}-\beta\Sigma^{-1}))m_{t}-P_{t}\beta\Sigma^{-1}m\Bigr]=-\frac{\gamma\beta}{\kappa}P_{t}\Sigma^{-1}\,(m_{t}-m),\end{aligned}$ (34)
$\kern 1.5pt\frac{\mathrm{d}}{\mathrm{d}t}\Sigma_{t}=-\frac{\gamma}{\kappa}(I-\frac{\beta}{\kappa}P_{t}\mathbf{C}_{*,t}^{-1})\Sigma_{t}-\frac{\gamma}{\kappa}\Sigma_{t}\,(I-\frac{\beta}{\kappa}\mathbf{C}_{*,t}^{-1}P_{t})+2\mathbf{C}_{*,t}.$ (35)
@@
    $\gamma=\Bigl(\frac{1}{\lambda}+\alpha+1\Bigr)^{-1}\kappa+\frac{\beta}{\beta+1}.$ (38)
@@
If $\pi$ is a stationary distribution of localized CBS, then $\partial_{t}\rho_{t}(u)=0$} when $\rho_{t}=\pi$}. Since then both $\rho_{t}$ and $\pi$ are Gaussian, we can use Lemma 4.1, fill in an expression for the preconditioner such as Eq. 32, and require $\frac{\mathrm{d}}{\mathrm{d}t}m_{t}=0$ and $\frac{\mathrm{d}}{\mathrm{d}t}\Sigma_{t}=0$ when $m_{t}=m$ and $\Sigma_{t}=\Sigma$. Clearly, $\frac{\mathrm{d}}{\mathrm{d}t}m_{t}$ as given by Eq. 31a is zero irrespective of $\gamma$ and $\mathbf{C}_{*,t}$ when $m_{t}=m$. We also have
$\frac{\mathrm{d}}{\mathrm{d}t}\Sigma_{t}$ $=2\Bigl[\mathbf{C}_{*,t}-\frac{\gamma}{\kappa}\Sigma+\frac{\gamma\beta}{2\kappa^{2}}\left(P_{t}\mathbf{C}_{*,t}^{-1}\Sigma+\Sigma\mathbf{C}_{*,t}^{-1}P_{t}\right)\Bigr].$ (39)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\Sigma_{t}$ $=2\Bigl[\mathbf{C}_{*,t}+\frac{\gamma}{\kappa}\Bigl(\bigl(\kappa\frac{\beta+1}{\beta}\mathbf{C}_{*,t}\Sigma^{-1}+I\bigr)^{-1}-I\Bigr)\Sigma\Bigr]$ (40)
@@
Equation 32d becomes $\mathbf{C}_{*,t}=(\lambda^{-1}+\alpha+1)\Sigma$ when $\Sigma_{t}=\Sigma$, which means that Eq. 41 holds when $\gamma$ is given by Eq. 38. The other cases correspond to specific choices for $\lambda$ and $\alpha$. ∎
@@
Remark 4.3 (Non-Gaussian performance).
In the non-Gaussian case, localized CBS with finite parameters $(\beta,\kappa)$ will not sample from the exact target distribution $\pi$ in general. Characterizing the well-posedness of Eq. 28 and the relation between $\pi$ and a possible stationary distribution $\rho_{\infty}$ of localized CBS is an open question.
@@ 4.4 Affine-invariance
The SDE Eq. 28 is affine-invariant (see Definition 2.1) when the preconditioner is given by $\mathbf{C}(\rho_{t};U_{t})=\mathcal{C}(\rho_{t})$, $\mathbf{C}(\rho_{t};U_{t})=\mathcal{C}_{\alpha}(\rho_{t})$, or $\mathbf{C}(\rho_{t};U_{t})=\mathcal{C}_{\alpha,\lambda\mathcal{C}(\rho_{t})}(\rho_{t};U_{t})$.
@@
$\partial_{t}\bar{\rho}_{t}(z)$ $=\mathinner{\!\left\lvert M\right\rvert}\partial_{t}\rho_{t}(u)$
$=\mathinner{\!\left\lvert M\right\rvert}\nabla_{u}\cdot\Bigl(\frac{\gamma}{\kappa}\rho_{t}(u)\,(u-\mu_{\beta,\kappa\mathbf{C}^{\widehat{\pi}}(\rho_{t};u)}^{\widehat{\pi}}(\rho_{t};u))+\mathbf{C}^{\widehat{\pi}}(\rho_{t};u)\nabla_{u}\rho_{t}(u)\Bigr)$
@@ 5 Comparison to alternative methods
This section compares the dynamics Eq. 28 of localized CBS to those of various other McKean–Vlasov-based samplers. Subsection 5.1 discusses classical CBS, after which subsections 5.2 and 5.3 compares our method to two algorithms that use localization in a similar way. Various other dynamics are briefly discussed in subsection 5.4. A detailed numerical comparison to some of these methods follows in section 7; here, we focus on different aspects of the dynamics and how these may affect the properties of the resulting algorithms.
@@ 5.1 Consensus-based sampling
Let $U_{t}\sim\mathcal{N}(m_{t},\Sigma_{t})$ follow the localized CBS dynamics Eq. 28 preconditioned with the weighted covariance $\mathbf{C}(\rho;u)=\mathcal{C}_{\alpha,\lambda\mathcal{C}(\rho)}(\rho;u)$ and let $U_{t}^{\prime}\sim\mathcal{N}(m_{t}^{\prime},\Sigma_{t}^{\prime})$ follow the CBS dynamics Eq. 15. Let both the target density $\pi=\mathcal{N}(m,\Sigma)$ and the initial densities of both dynamics $\mathcal{N}(m_{0},\Sigma_{0})=\mathcal{N}(m_{0}^{\prime},\Sigma_{0}^{\prime})$} be Gaussian.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\Sigma^{\prime}_{t}=-2\alpha^{\prime}\Sigma^{\prime}_{t}\,(\Sigma+\alpha^{\prime}\Sigma^{\prime}_{t})^{-1}(\Sigma^{\prime}_{t}-\Sigma).$ (49)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\Sigma_{t}$ $=2\Bigl[\mathbf{C}-2\frac{\alpha^{\prime}}{\alpha(\alpha^{\prime}+1)}\Sigma_{t}+\frac{(\alpha^{\prime})^{2}}{\alpha^{2}(\alpha^{\prime}+1)}\left(P_{t}\mathbf{C}^{-1}\Sigma_{t}+\Sigma_{t}\mathbf{C}^{-1}P_{t}\right)\Bigr]$
@@
where we noted that $\gamma=2\frac{\alpha^{\prime}}{\alpha^{\prime}+1}$. This is Eq. 49 accelerated with a factor $c$; hence, $\Sigma_{t}=\Sigma^{\prime}_{ct}$.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}m^{\prime}_{t}=-\alpha^{\prime}\Sigma^{\prime}_{t}\,(\Sigma+\alpha^{\prime}\Sigma^{\prime}_{t})^{-1}(m^{\prime}_{t}-m).$ (50)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}m_{t}$ $=-2\alpha^{\prime}cP_{t}\Sigma^{-1}\,(m_{t}-m)=-c\alpha^{\prime}\Sigma_{t}\,(\Sigma+\alpha^{\prime}\Sigma_{t})^{-1}(m_{t}-m)$